OpenStudy (anonymous):

Whats the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6.

OpenStudy (anonymous):

slope is 3 so use \[y+2=3(x-3)\] as a start

OpenStudy (aroub):

parallel=same slope y=mx+b , m=3 , (3,-2) -2=3(3)+b -2=9+b -9-2=b -11=b y=3x-11